# and quantum information science

### v2

@@ 5.4 Physical criteria for complexity growth: general results and applications
At least at the early stages of time evolution, we can be sure from the results discussed above that the trajectory generated by the Hamiltonian is the shortest between $\mathbbm{1}$ and $e^{-iHt}$. How could this linear geodesic trajectory stop being the shortest? One possibility is the existence of a geodesic loop that allows us to exploit the topology of $\mathrm{SU}(N)$ to arrive at the same destination through a different path. This is what happened in the $\mathrm{SU}(2)$ example above. Geodesic loops are difficult to work out in the general unitary group, so we will not attempt to do that. However, note that in $\mathrm{SU}(2)$ the shorter path arising from a great cycle on the group manifold became relevant after the time evolution trajectory had traversed a diameter of the manifold. We expect something like that to be also generally true for geodesic loops of $\mathrm{SU}(N>2)$. Following Ref. [19], we also expect the diameter of $\mathrm{SU}(N)$ to be exponentially large in the number of degrees of freedom in complexity metrics satisfying reasonable consistency criteria. This means that, all else being equal, we expect the geodesic loop obstruction to start to play a role in complexity growth after exponential time. So we will focus on another kind of shortcut on the group manifold that could truncate the linear complexity growth in (33) far earlier. These shortcuts are associated to conjugate points on the linear Hamiltonian trajectory that we will now describe.
A conjugate point occurs on a geodesic $U(s):[0,1]\to\mathrm{SU}(N)$ with $U(0)=P$ and $U(1)=Q$ if a perturbation of the path produces a new curve with the same boundary conditions that satisfies the geodesic equation to first order. This implies that the geodesic we started with is a saddlepoint, not a minimum of the length, so that we can find a shorter path. The actual minimizing geodesic signaled by the appearance of conjugate point can be a finite distance away. Since conjugate points indicate converging geodesics, they cannot occur if all the sectional curvatures are negative. There is a result of Milnor [119] that says that most sectional curvatures of Lie groups are negative. So some authors have considered high genus Riemann surfaces with a metric induced from the covering space, a hyperbolic disk, as a toy model of the unitary group manifold [50]. In such a toy model there are no conjugate points, so they are irrelevant. Others have constructed special metrics on the unitary group that are designed to push the conjugate points to infinity [97]. But these approaches seem misleading about the structure of the unitary group because the main result of Milnor is that any right invariant metric on $\mathrm{SU}(N)$ with $N>2$ must have some positive sectional curvature, or else it is flat. Indeed, as we will see, conjugate points are some of the earliest obstructions to complexity growth in free and integrable theories.
@@
$i\frac{d\delta V_{L}(s)}{ds}=\mu t[H,\delta V_{NL}(s)]_{L}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ i\frac{d\delta V_{NL}(s)}{ds}=\frac{\mu t}{1+\mu}+[H,\delta V_{NL}(s)]_{NL}$ (34)
@@
By solving this equation, we can work out how an initial perturbation of the trajectory at the origin changes the trajectory. The authors of [112] used the formal solution of the above equation to write down a superoperator $Y[\delta V(0)]$ mapping the initial perturbation of the trajectory, $\delta V(0)$, onto a perturbation of the endpoint. Conjugate points occur when the endpoint is unchanged, namely when the superoperator $Y$ has a zero mode in the formalism of [112]. We do not write the superoperator explicitly here because it has a complicated form that requires further definitions to explain, and instead refer the reader to [112]. By working out criteria for the appearance of such zero modes, the authors of [112, 114] arrived at a number of general results for the appearance of conjugate points that are described below:
@@ 1. Result 1:
    For arbitrary local Hamiltonians with a finite complexity cost factor $\mu$, conjugate points must exist at finite distance along the linear geodesic. This is proved by starting with the result of [120] that conjugate points exist at finite distance when the complexity penalty $\mu$ vanishes. The persistence of the conjugate points is then established by a continuity argument as $\mu$ increases, along with Morse theory on the space of paths on the group [114]. So the absence of conjugate points must be shown to establish linear complexity growth until times of order $t\sim e^{S}$.
@@ 2. Result 2:
    If a q-local Hamiltonian has an adjoint eigenoperator $O$ such that $[H,O]=\lambda O$ with $\lambda\in\mathbb{R}$ and $O\in\{{\rm local/easy\ operators}\}$, then conjugate points occur at times $t_{*}=\frac{2\pi}{\lambda}\mathbb{Z}$ (where $\mathbb{R}$ are the reals and $\mathbb{Z}$ are the integers). This is proven in [114] by explicit examination of the zero mode condition on the superoperator $Y$ mentioned above.
@@ 3. Result 3:
    If a q-local Hamiltonian has an adjoint eigenoperator $O^{\prime}$ such that $[H,O^{\prime}]=\lambda^{\prime}O^{\prime}$ with $\lambda^{\prime}\in\mathbb{R}$ and $O^{\prime}\in\{{\rm nonlocal/hard\ operators}\}$, then conjugate points occur at times $t_{*}=\frac{2\pi(1+\mu)}{\lambda^{\prime}}\mathbb{Z}$. This is proven in the same way as Result 2 above.
@@ 4. Result 4:
[…]
@@ 5. Result 5:
    where $r_{mn}=\mathcal{O}(1)$ for $m\neq n$ and ${\rm poly}(S)$ means polynomial in $S$. This criterion, which is called the Eigenstate Complexity Hypothesis (ECH) (since it recalls the Eigenstate Thermalization Hypothesis), says that energy eigenstates cannot be mapped onto each other by easy operators. This result was proven in [112] by showing that if we assume ECH, the zero modes of the superoperator $Y$ occur at times $t^{*}=(2\pi/\Delta_{{\rm max}})(1+e^{\epsilon S})$, where $\Delta_{\rm max}$ is the largest gap in the spectrum.
These theorems can be applied to the SYK model with $n$ Majorana fermions, which we will return to in Sections 6.3.3 and 7.5.1. Here we briefly summarize the results of [112, 114] in this regard. The authors provide evidence that: (a) The Nielsen complexity of the free SYK model (with only quadratic interactions) is bounded as $\mathcal{C}(t)\lesssim\sqrt{n}$; (b) The Nielsen complexity of integrable deformations of the free SYK model is bounded as $\mathcal{C}(t)\lesssim{\rm poly}(n)$, meaning that the initial linear growth is guaranteed to truncate in polynomial time; (c) 35 above, and numerical analysis of the spectrum and eigenstates of chaotic deformations of the free SYK model suggest that the Nielsen complexity of such systems is bounded as $\mathcal{C}(t)\lesssim e^{\epsilon n}$, meaning that linear complexity growth can persist for exponential time, if we neglect possible global obstructions from geodesic loops. The finding (c) applies to any theory in which the energy eigenstates are a Haar random unitary rotation of a “standard” or “free” basis of states. The ETH hypothesis for chaotic theories suggests that their eigenstates may indeed be of this form.
@@ 5.5 Relation between Nielsen complexity and entanglement: binding complexity
The limiting case, where only operations acting between subsystems have non-vanishing cost, was referred to as binding complexity in [111]. There, the authors suggested that binding complexity quantifies the difficulty of distributing entanglement among multiple parties. The authors of Ref. [111] further suggested that the notion of binding complexity could serve as a measure for the robustness of entanglement. Here, one typically has in mind the $n$-party (with $n\geq 2$) generalizations of the W and GHZ states. The GHZ state is separable upon tracing out any subset of the parties, but the W state is not. In this sense, W states possess more robust entanglement. The authors of [111] argued that the binding complexity should scale as $\mathcal{O}(n)$ for GHZ states, and as $\mathcal{O}(n^{2})$ for $W$ states. The paper also presented results for the binding complexity of Gaussian states of free QFT. As we review next, we can get exact results for the binding complexity of spin chains in Nielsen’s framework when the cost of generators acting within each subsystem is small enough [115]. It turns out that cost functions satisfying this property relate Nielsen’s complexity to the Rényi min-entropy in a precise way. Moreover, the above cost functions are bounded by the entanglement entropy of a state over a subregion [121, 115].
To illustrate these results, let us begin with a formal definition of the binding complexity. Consider splitting a quantum system (for instance, a spin chain) into $m$ subsystems $A_{k=1,2,\ldots,m}\,.$ The binding complexity is the minimal number of gates, each acting on a limited number of subsystems, needed to implement a target unitary with cost only incurred for gates acting between subsystems [111]. Below, we focus on 2-local gates which act on, at most, two subsystems each. Within Nielsen’s framework, binding complexity has a close analogue, obtained by choosing a particular cost function on the unitary space. To illustrate, consider a bipartite system whose Hamiltonian (17) decomposes as [^15] We denote the generators of the special unitary group in subsystem $A_{1}$ with index $a\in\{1,\dots,N_{A_{1}}^{2}-1\}$ from the beginning of the alphabet, and the generators in $A_{2}$ with index $i\in\{1,\dots,N_{A_{2}}^{2}-1\}$ from the middle part of the alphabet. We also denote with a sub(super)script $A_{1},A_{2}$ the subsystem where the local generators act.
@@
We can analytically compute the binding state complexity for certain choices of the penalty factors $q_{ai}$ in Eq. (37), and of the parameter $p$ in Eq. (25). A Riemannian cost function is our first example. Let us write out the binding state complexity associated with the cost function (25), $p=2$, and the penalty factors (37) with $q_{ai}=1$ [115]:
@@
The hom stands for the nonlocal generators’ homogeneous penalties. The $\mathcal{N}_{A_{k}}$ are normalization factors defined such that $\mathop{\rm tr}(T^{A_{k}}_{a}T^{A_{k}}_{b})=\mathcal{N}_{A_{k}}\delta_{ab}$. The $S_{\infty}$ denotes the Rényi min-entropy of the reduced state of subsystem $A_{1}$ in the target state $\lvert\psi_{\rm T}\rangle$.
@@
The relationship between binding complexity and Schmidt coefficients leads not only to exact results, but also bounds other quantum-information quantities. For instance, one can lower-bound the binding state complexity, in terms of the entanglement entropy of a state over a subregion $A_{1}$ [122, 123, 124, 121]. Denote by $S_{A_{1}}$ the entanglement entropy of subregion $A_{1}$, by $d$ the dimensionality of the smaller subsystem’s Hilbert space, and by $c$ a constant. The $p{=}1$ binding state complexity upper bounds $S_{A_{1}}$ as follows:
@@
Binding complexity also provides lower bounds for other notions of complexity, such as the geometrically local complexity [121]. Further details of the latter quantity will be discussed around Eq. (156).
Finally, let us mention that in the context of the holographic correspondence, binding complexity has been conjectured to be dual to the interior volume of multi-boundary wormholes [111].
@@ 5.6 Complexity in quantum field theories
As we will review in Sec. 7, Nielsen’s complexity experienced a revival following a conjectured relationship with gravitational quantities in the context of the holographic AdS/CFT correspondence. The holographic correspondence deals with Quantum Field Theories (QFTs) – many-body quantum systems which do not conserve particle number. To allow a quantitative comparison with holographic conjectures, it was therefore necessary to develop a method for calculating complexity in QFTs. Naturally, the first examples studied were those of free QFTs, first bosonic [99, 100] and then fermionic [103, 104], as reviewed in subsection 5.6.1. While these calculations are too simple for an exact comparison with chaotic theories in general – and with holographyic systems in particular – this approach helped to bridge the gap between heuristic arguments and more rigorous calculations. This framework also inspired later studies of complexity in more complicated field theories, including weakly interacting theories [106, 125], and conformal field theories, as described in subsections 5.6.2 and 5.6.3 below.

### v3

@@ 5.4 Physical criteria for complexity growth: general results and applications
At least at the early stages of time evolution, we can be sure from the results discussed above that the trajectory generated by the Hamiltonian is the shortest between $\mathbbm{1}$ and $e^{-iHt}$. How could this linear geodesic trajectory stop being the shortest? One possibility is the existence of a geodesic loop that allows us to exploit the topology of $\mathrm{SU}(N)$ to arrive at the same destination through a different path. This is what happened in the $\mathrm{SU}(2)$ example above. Geodesic loops are difficult to work out in the general unitary group, so we will not attempt to do that. However, note that in $\mathrm{SU}(2)$ the shorter path arising from a great cycle on the group manifold became relevant after the time evolution trajectory had traversed a diameter of the manifold. We expect something like that to be also generally true for geodesic loops of $\mathrm{SU}(N>2)$. Following Ref. Brown:2022phc, we also expect the diameter of $\mathrm{SU}(N)$ to be exponentially large in the number of degrees of freedom in complexity metrics satisfying reasonable consistency criteria. This means that, all else being equal, we expect the geodesic loop obstruction to start to play a role in complexity growth after exponential time. So we will focus on another kind of shortcut on the group manifold that could truncate the linear complexity growth in (33) far earlier. These shortcuts are associated to conjugate points on the linear Hamiltonian trajectory that we will now describe.
A conjugate point occurs on a geodesic $U(s):[0,1]\to\mathrm{SU}(N)$ with $U(0)=P$ and $U(1)=Q$ if a perturbation of the path produces a new curve with the same boundary conditions that satisfies the geodesic equation to first order. This implies that the geodesic we started with is a saddlepoint, not a minimum of the length, so that we can find a shorter path. The actual minimizing geodesic signaled by the appearance of conjugate point can be a finite distance away. Since conjugate points indicate converging geodesics, they cannot occur if all the sectional curvatures are negative. There is a result of Milnor Milnor:1976pu that says that most sectional curvatures of Lie groups are negative. So some authors have considered high genus Riemann surfaces with a metric induced from the covering space, a hyperbolic disk, as a toy model of the unitary group manifold Brown:2016wib. In such a toy model there are no conjugate points, so they are irrelevant. Others have constructed special metrics on the unitary group that are designed to push the conjugate points to infinity Brown:2021uov. But these approaches seem misleading about the structure of the unitary group because the main result of Milnor is that any right invariant metric on $\mathrm{SU}(N)$ with $N>2$ must have some positive sectional curvature, or else it is flat. Indeed, as we will see, conjugate points are some of the earliest obstructions to complexity growth in free and integrable theories.
@@
$i\frac{d\delta V_{L}(s)}{ds}=\mu t[H,\delta V_{NL}(s)]_{L}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ i\frac{d\delta V_{NL}(s)}{ds}=\frac{\mu t}{1+\mu}+[H,\delta V_{NL}(s)]_{NL}$ (34)
@@
By solving this equation, we can work out how an initial perturbation of the trajectory at the origin changes the trajectory. The authors of Balasubramanian:2019wgd used the formal solution of the above equation to write down a superoperator $Y[\delta V(0)]$ mapping the initial perturbation of the trajectory, $\delta V(0)$, onto a perturbation of the endpoint. Conjugate points occur when the endpoint is unchanged, namely when the superoperator $Y$ has a zero mode in the formalism of Balasubramanian:2019wgd. We do not write the superoperator explicitly here because it has a complicated form that requires further definitions to explain, and instead refer the reader to Balasubramanian:2019wgd. By working out criteria for the appearance of such zero modes, the authors of Balasubramanian:2019wgd, Balasubramanian:2021mxo arrived at a number of general results for the appearance of conjugate points that are described below:
@@ 1. Result 1:
    For arbitrary local Hamiltonians with a finite complexity cost factor $\mu$, conjugate points must exist at finite distance along the linear geodesic. This is proved by starting with the result of Dowling:2006tnk that conjugate points exist at finite distance when the complexity penalty $\mu$ vanishes. The persistence of the conjugate points is then established by a continuity argument as $\mu$ increases, along with Morse theory on the space of paths on the group Balasubramanian:2021mxo. So the absence of conjugate points must be shown to establish linear complexity growth until times of order $t\sim e^{S}$.
@@ 2. Result 2:
    If a q-local Hamiltonian has an adjoint eigenoperator $O$ such that $H,O]=\lambda O$ with $\lambda\in\mathbb{R}$ and $O\in\{{\rm local/easy\ operators}\}$, then conjugate points occur at times $t_{*}=\frac{2\pi}{\lambda}\mathbb{Z}$ (where $\mathbb{R}$ are the reals and $\mathbb{Z}$ are the integers). This is proven in Balasubramanian:2021mxo by explicit examination of the zero mode condition on the superoperator $Y$ mentioned above.
@@ 3. Result 3:
    If a q-local Hamiltonian has an adjoint eigenoperator $O^{\prime}$ such that $H,O^{\prime}]=\lambda^{\prime}O^{\prime}$ with $\lambda^{\prime}\in\mathbb{R}$ and $O^{\prime}\in\{{\rm nonlocal/hard\ operators}\}$, then conjugate points occur at times $t_{*}=\frac{2\pi(1+\mu)}{\lambda^{\prime}}\mathbb{Z}$. This is proven in the same way as Result 2 above.
@@ 4. Result 4:
    Let $M_{\alpha\beta}(t)=\int_{0}^{1}ds\int_{0}^{1}ds^{\prime}\,{\rm Tr}[e^{i(s-s^{\prime})tH}T_{\alpha}e^{-i(s-s^{\prime})tH}T_{\beta}]$, where $T_{\alpha}$ and $T_{\beta}$ are any two local/easy operators. Suppose $M_{\alpha\beta}(t)$ has an eigenvector $X^{\alpha}$, with vanishing eigenvalue at some time $t^{*}$, then a conjugate point occurs at $t^{*}$. We can think of $M_{\alpha\beta}$ as a temporally smeared, infinite temperature 2-point function of local operators. To prove this, given the zero mode $X^{\alpha}$, the authors of Balasubramanian:2021mxo expanded an initial perturbation of the linear geodesic as $\delta V(0)=\sum_{\alpha}X^{\alpha}T_{\alpha}$ where $T_{\alpha}$ are the local/easy generators, and showed that the Frobenius norm of the superoperator $Y$ discussed above vanishes when it acts on this $\delta V(0)$. This implies that the superoperator vanishes on this perturbation, which by construction implies a conjugate point.
@@ 5. Result 5:
    where $r_{mn}=\mathcal{O}(1)$ for $m\neq n$ and ${\rm poly}(S)$ means polynomial in $S$. This criterion, which is called the Eigenstate Complexity Hypothesis (ECH) (since it recalls the Eigenstate Thermalization Hypothesis), says that energy eigenstates cannot be mapped onto each other by easy operators. This result was proven in Balasubramanian:2019wgd by showing that if we assume ECH, the zero modes of the superoperator $Y$ occur at times $t^{*}=(2\pi/\Delta_{{\rm max}})(1+e^{\epsilon S})$, where $\Delta_{\rm max}$ is the largest gap in the spectrum.
These theorems can be applied to the SYK model with $n$ Majorana fermions, which we will return to in Sections 6.3.3 and 7.5.1. Here we briefly summarize the results of Balasubramanian:2019wgd, Balasubramanian:2021mxo in this regard. The authors provide evidence that: (a) The Nielsen complexity of the free SYK model (with only quadratic interactions) is bounded as $\mathcal{C}(t)\lesssim\sqrt{n}$; (b) The Nielsen complexity of integrable deformations of the free SYK model is bounded as $\mathcal{C}(t)\lesssim{\rm poly}(n)$, meaning that the initial linear growth is guaranteed to truncate in polynomial time; (c) 35 above, and numerical analysis of the spectrum and eigenstates of chaotic deformations of the free SYK model suggest that the Nielsen complexity of such systems is bounded as $\mathcal{C}(t)\lesssim e^{\epsilon n}$, meaning that linear complexity growth can persist for exponential time, if we neglect possible global obstructions from geodesic loops. The finding (c) applies to any theory in which the energy eigenstates are a Haar random unitary rotation of a “standard” or “free” basis of states. The ETH hypothesis for chaotic theories suggests that their eigenstates may indeed be of this form.
@@ 5.5 Relation between Nielsen complexity and entanglement: binding complexity
The limiting case, where only operations acting between subsystems have non-vanishing cost, was referred to as binding complexity in Balasubramanian:2018hsu. There, the authors suggested that binding complexity quantifies the difficulty of distributing entanglement among multiple parties. The authors of Ref. Balasubramanian:2018hsu further suggested that the notion of binding complexity could serve as a measure for the robustness of entanglement. Here, one typically has in mind the $n$-party (with $n\geq 2$) generalizations of the W and GHZ states. The GHZ state is separable upon tracing out any subset of the parties, but the W state is not. In this sense, W states possess more robust entanglement. The authors of Balasubramanian:2018hsu argued that the binding complexity should scale as $\mathcal{O}(n)$ for GHZ states, and as $\mathcal{O}(n^{2})$ for $W$ states. The paper also presented results for the binding complexity of Gaussian states of free QFT. As we review next, we can get exact results for the binding complexity of spin chains in Nielsen’s framework when the cost of generators acting within each subsystem is small enough Baiguera:2023bhm. It turns out that cost functions satisfying this property relate Nielsen’s complexity to the Rényi min-entropy in a precise way. Moreover, the above cost functions are bounded by the entanglement entropy of a state over a subregion eisert2021entangling, Baiguera:2023bhm.
To illustrate these results, let us begin with a formal definition of the binding complexity. Consider splitting a quantum system (for instance, a spin chain) into $m$ subsystems $A_{k=1,2,\ldots,m}\,.$ The binding complexity is the minimal number of gates, each acting on a limited number of subsystems, needed to implement a target unitary with cost only incurred for gates acting between subsystems Balasubramanian:2018hsu. Below, we focus on 2-local gates which act on, at most, two subsystems each. Within Nielsen’s framework, binding complexity has a close analogue, obtained by choosing a particular cost function on the unitary space. To illustrate, consider a bipartite system whose Hamiltonian (17) decomposes as [^17] We denote the generators of the special unitary group in subsystem $A_{1}$ with index $a\in\{1,\dots,N_{A_{1}}^{2}-1\}$ from the beginning of the alphabet, and the generators in $A_{2}$ with index $i\in\{1,\dots,N_{A_{2}}^{2}-1\}$ from the middle part of the alphabet. We also denote with a sub(super)script $A_{1},A_{2}$ the subsystem where the local generators act.
@@
We can analytically compute the binding state complexity for certain choices of the penalty factors $q_{ai}$ in Eq. (37), and of the parameter $p$ in Eq. (25). A Riemannian cost function is our first example. Let us write out the binding state complexity associated with the cost function (25), $p=2$, and the penalty factors (37) with $q_{ai}=1$ Baiguera:2023bhm:
@@
The hom stands for the nonlocal generators’ homogeneous penalties. The $\mathcal{N}_{A_{k}}$ are normalization factors defined such that $\mathop{\rm tr}(T^{A_{k}}_{a}T^{A_{k}}_{b})=\mathcal{N}_{A_{k}}\delta_{ab}$. The $S_{\infty}$ denotes the Rényi min-entropy of the reduced state of subsystem $A_{1}$ in the target state $\lvert\psi_{\rm T}\rangle$. [^18] The Rényi min-entropy of the reduced state of a subsystem $A_{1}$ in the state $|\psi_{\rm T}\rangle$ is defined by $S_{\infty}(|\psi_{\rm T}\rangle)=-2\log\bar{\lambda}_{1}$, where $\bar{\lambda}_{1}$ is the maximal Schmidt coefficient of the target state with respect to the same bipartite division.
@@
The relationship between binding complexity and Schmidt coefficients leads not only to exact results, but also bounds other quantum-information quantities. For instance, one can lower-bound the binding state complexity, in terms of the entanglement entropy of a state over a subregion $A_{1}$ bravyi:2007, van:2013, Verstraete:2016, eisert2021entangling. Denote by $S_{A_{1}}$ the entanglement entropy of subregion $A_{1}$, by $d$ the dimensionality of the smaller subsystem’s Hilbert space, and by $c$ a constant. The p=1p{=}1 binding state complexity upper bounds $S_{A_{1}}$ as follows:
@@
Binding complexity also provides lower bounds for other notions of complexity, such as the geometrically local complexity eisert2021entangling. Further details of the latter quantity will be discussed around Eq. (156).
Finally, let us mention that in the context of the holographic correspondence, binding complexity has been conjectured to be dual to the interior volume of multi-boundary wormholes Balasubramanian:2018hsu.
@@ 5.6 Complexity in quantum field theories
As we will review in Sec. 7, Nielsen’s complexity experienced a revival following a conjectured relationship with gravitational quantities in the context of the holographic AdS/CFT correspondence. The holographic correspondence deals with Quantum Field Theories (QFTs) – a description of many-body quantum systems in terms of fields. To allow a quantitative comparison with holographic conjectures, it was therefore necessary to develop a method for calculating complexity in QFTs. Naturally, the first examples studied were those of free QFTs, first bosonic Jefferson:2017sdb, Chapman:2017rqy and then fermionic Khan:2018rzm, Hackl:2018ptj, as reviewed in subsection 5.6.1. While these calculations are too simple for an exact comparison with chaotic theories in general – and with holographyic systems in particular – this approach helped to bridge the gap between heuristic arguments and more rigorous calculations. This framework also inspired later studies of complexity in more complicated field theories, including weakly interacting theories Bhattacharyya:2018bbv, Cotler:2018ufx, and conformal field theories, as described in subsections 5.6.2 and 5.6.3 below.